# Context versus Prior Knowledge in Language Models

### v1

@@ 4.2.2 Validating Susceptibility Scores
We compare susceptibility scores to other metrics as further evidence for the meaningfulness of our metric. Since we are, to our knowledge, the first to conceptualize per-entity scores for susceptibility, we adapt the memorization ratio (MR) metric from Longpre et al. (2021) to apply on a per-entity basis, which is described in more detail in App. D. We compare each entity's susceptibility score and entity-specific MR for each query, using the same setup as in § 4.1 for the Pythia-6.9b-deduped model.
In Fig. 3, we note a decreasing upper bound relationship between susceptibility scores and MR for both open queries and closed queries. In order to interpret these results, it is important to note that these metrics differ in the nature of what they measure—susceptibility measures how the entropy of the output distribution is affected by context, whereas MR counts the actual outputs sampled from the model. Depending on the model's confidence without the prepended context, it is possible that both large and small changes in entropy (corresponding to large and small susceptibility scores) could indicate a change of output (low MR), while an unchanged output (high MR) indicates that the entropy likely remained similar (low susceptibility). Thus, we do not expect to see a straightforward correlation between the two metrics and instead expect a negative correlation between the upper bound of the susceptibility score at a given entity-specific MR. Both open and closed queries appear to exhibit this relationship (at least at the outliers), supporting the convergent validity of susceptibility. The open queries further show a decreasing pattern at each quartile for all bins except the lowest one (MR between 0 and 0.2). The different behavior for the lowest bin might be explained by a limitation of MR; especially for open queries, a model may fail to produce the original answer for most or all contexts if it does not know it from the start. In this case, MR cannot distinguish between the model behavior for these difficult queries, whereas susceptibility scores can provide more granular information about model behavior for such entities. This effect might be exacerbated by the majority of entities falling into the bin with the lowest MR for both open and closed queries. Additionally, the closed queries appear to have a mostly increasing pattern at each quartile across the bins. This could be explained by an artifact of the construction of the closed queries. Since the original answer of all closed queries is Yes, it is possible that the contexts increase the confidence in Yes due to some model artifact or token bias, which would explain higher susceptibility scores even for higher MR.
@@
Following the same setup used in § 4.2.1, we consider three seeds for sampling contexts and the same two query forms for both open and closed questions to test for reliability (low variance) for susceptibility scores across both axes of seeds/sampled contexts and query forms. Fig. 4 shows strong evidence for the reliability of susceptibility scores. Similarly to the persuasion scores, the variance for both open and closed queries is very low when taken across seeds and closed query forms; open query forms have higher variance because they have meaningfully different possible answers, as discussed in § 4.2.1.
@@ 4.3.1 Relevance
We use the setup described in § 4.1. For the relevance test, we consider a relevant context to be one that mentions the queried entity, and an irrelevant context as one that does not. We hypothesize that relevant contexts should be more persuasive than irrelevant ones. For each of the 122 queries, we find the mean persuasion score of the relevant contexts and irrelevant contexts, and use a permutation test to determine whether the mean persuasion score for relevant contexts (across entities) is higher than the mean persuasion score for irrelevant contexts, using a significance level of with the BH correction.
@@ Results.
As seen in Fig. 5 (), across most model sizes and relations, relevant contexts are significantly more persuasive than irrelevant contexts. Specifically, depending on the model, 95–100% of open queries and 83–100% of closed queries showed a significant result. We also see a trend in which, as model size increases, so too does the degree to which relevant contexts are more persuasive (as measured by the mean effect size). We summarize the significance test results and effect sizes for all models and queries in App. E.
@@ 4.3.2 Assertiveness
We use the same setup as in § 4.3.1 but instead explore our hypothesis that assertive contexts are more persuasive than the base ones by computing the mean persuasion score of each group and testing for a significant difference for each query.
@@ Results.
Fig. 5 () shows that, consistently across model sizes, assertive contexts tend to be more persuasive than base contexts for closed queries but not for open queries. Moreover, assertive contexts are most persuasive for medium-sized models such as 1.4b and 2.8b. We hypothesize that smaller models may be less persuaded by assertive contexts because they may be worse at integrating context into its answer. Larger models may be less persuaded because of their stronger prior knowledge/lower susceptibility to context, whether assertive or not. We highlight the significance test results and effect sizes for all queries for the Pythia-6.9b-deduped model in Fig. 6 and other models in App. E.
@@
The permutation tests suggest evidence for a significant difference in 88% of the closed queries and 74% of the open queries for the Pythia-6.9b-deduped model. However, there is not a consistent directional pattern; from Fig. 7, we see that negations are significantly more persuasive for some queries while significantly less persuasive for others. This pattern is similar for other models as seen in App. E. Fig. 5 () also shows that the smallest model is the most sensitive to being persuaded differently by negation vs base contexts. We hypothesize this may be due to potential spurious correlations/token biases between, for example, seeing the word not and the model's probability of outputting No.
@@ 4.3.4 Comparing Context Qualities
For all models on open queries, the relevance of a context has the greatest effect on persuasion score compared to assertiveness or negation (as measured by effect size). This is consistent with our intuition that the model should be more sensitive to whether the queried entity is mentioned in the context than to other context features. However, we can see in Fig. 5 that for medium-sized models on closed queries, the other context comparisons have stronger effect sizes (e.g., assertiveness and negation for the 2.8b model). This could be explained by potential spurious correlations/token biases associating, i.e., the word not with the model outputting No or the word definitely with outputting Yes.
@@ 4.4.1 Familiar vs Unfamiliar Entities
We use the setup described in § 4.1 to compute susceptibility scores for known and unknown entities. We use a permutation test with and the BH correction to test for whether the known real entities are less susceptible than the unknown fake entities.101010We filter out fake entities that appear in the training data to better represent unknown entities in our analysis. The detailed setup can be found in App. C.
@@ Results.
For the Pythia-6.9b-deduped model, we find that with queries (open questions) and queries (closed questions), familiar entities have significantly lower susceptibility scores than unfamiliar fake entities. We conjecture that for the remaining queries, the model may not have strong prior biases about the sampled entities for these queries. Indeed, further analysis (§ F.1) finds some evidence supporting the hypothesis that queries with smaller effect sizes/less significant -values feature less familiar entities, as we find a small correlation between effect size and entity frequency in the training set. Fig. 8 shows the distribution of susceptibility scores for real and fake entities for an example query with a particularly strong effect size.
@@ 4.4.2 Degrees of Familiarity
Since language models are parameterized with knowledge from their training corpora, we hypothesize that the model is less susceptible for entities with which it is more familiar, i.e., more frequently occurring in the training data. We investigate this relationship between the Pythia models behavior and frequency statistics in the Pile dataset on which they were trained Gao et al. (2020). In an attempt to capture the model's familiarity with the relation between the entity and answer, we count the number of co-occurrences between the entity and its corresponding answer within a 50-word window. We compare the susceptibility score to this co-occurrence frequency and find a significant correlation (Spearman ) for the Pythia-6.9b-deduped model. We see in Fig. 9 that as the training data frequency increases, the susceptibility scores' upper bound decreases. This trend is shared across all model sizes (see § F.2).
@@
Training data can be noisy, difficult, and expensive to search through for entity co-occurrences and more complex frequency statistics. Knowledge graphs offer a more precise alternative as they represent entities and relations extracted from common corpora like Wikipedia in a structured way. For example, within the pretraining data, it is very difficult to identify the number of different answers with which an entity will co-occur within the context of a specific relation. However, with a knowledge graph, we can easily identify the exact number of objects with which an entity might share a given relation. Thus, we explore the relationship between the relation-dependent degree of an entity in a knowledge graph and the susceptibility score. Similarly to the comparison against training data frequency, Fig. 10 also appears to show a decreasing upper bound between susceptibility scores and the degree for the Pythia-6.9b-deduped model; § F.2 further shows this trend is shared across all model sizes. We also provide the detailed methodology in § F.2. Such a trend suggests that unfamiliar entities have the potential to be highly susceptible, while very familiar entities tend to have lower susceptibility.
@@ 5 Applications
Social scientists use large language models (LLMs) for annotating data and descriptive data analysis, yet such use may inadvertently incorporate entity biases and skew model behavior (Ziems et al., 2024; Gilardi et al., 2023; Zhang et al., 2023; O'Hagan and Schein, 2023). To better contextualize model annotations for a case study on friend-enemy stance detection (Choi et al., 2016; Stoehr et al., 2023), we aim to understand how susceptibility scores may differ between friend and enemy-based entity pairs for the query The relationship between {entity1} and {entity2} is, e.g., are famous friend-based relationships more susceptible than enemy-based relationships? From Fig. 11, we see that with the Pythia-6.9b-deduped model for two specific query forms, enemy duos are less susceptible than friend duos, which can inform social scientists that a model's annotation for friend pairs may be more easily influenced by the context than enemy pairs. We provide more details in App. G.
@@
Since higher susceptibility scores indicate weaker induced biases for entities, we conjecture that this can relate to being underrepresented in the training data. Based on this, we consider how the susceptibility score can be used to study gender bias in LLMs. Using GPT-4, we collect stereotypically-biased contexts, gendered names, and neutral queries and run several experiments to identify gender discrepancies in susceptibility scores (see App. G for full details). We highlight a result where masculine names have a higher susceptibility than feminine names when swapping the genders in the stereotypical contexts, as seen in Fig. 12. This could indicate that the model is more surprised to see contexts claiming men follow feminine stereotypes, and therefore could suggest less representation of feminine stereotypes in the training data than masculine ones.
@@ 6 Discussion and Conclusion
We have made the case that the persuasion score and susceptibility score are both valid and reliable in measuring their respective constructs, following a well-established measurement modeling framework. Throughout our experiments, we find a common theme in the results: there is a strong, negative, possibly linear relationship between the upper bound of the susceptibility score and (a) the entity's memorization ratio (Fig. 3), (b) the log-co-occurrence frequency in the training data (Fig. 9), and (c) the log-relation-dependent degree in a knowledge graph (Fig. 10). That is, as each of those three values increases, we see a clear pattern indicating that the highest susceptibility scores tend to decrease. This is consistent with our hypothesis that the induced bias of an entity increases for a model as the model's expected familiarity with the entity increases. Furthermore, we find a difference in behavior between scores for open and closed queries; while in many experiments, we see similar patterns between the two, susceptibility and persuasion appear to have a stronger relationship with memorization ratio for open queries, while assertive contexts appear to be significantly more persuasive than base contexts primarily for closed queries. This difference is a surprising phenomenon that warrants future study; we hypothesize closed queries may behave this way due to common token biases (with its output space of Yes and No) and/or the influence of mentioning both an entity and an answer in the query. Finally, while we applied these metrics to analyze model behavior in two case studies, in future work we aim to apply them to unearth new perspectives in other context and prior knowledge-dependent problems such as retrieval-augmented generation, model editing and control, and few-shot learning.
@@
We face some technical limitations in executing the empirical aspects of this work. First, while § 3 defines the output space of as the set of all possible outputs , in practice, it is computationally expensive to estimate that probability distribution. Instead, we look only at the model's probability distribution of the next token, which could be a noisy signal, especially in cases where the answer suggested by a context and the answer suggested from prior knowledge share the same first token. Second, it is difficult to go through the whole Pile to count answer-entity co-occurrences without noise. Third, the scores depend on the sampled contexts, which may not be representative of all applications.
@@ Appendix A A Primer on Half-pointwise Mutual Information
where is the cross-entropy between and .
@@ Appendix B An Entity-Independent Persuasion Score
In addition to the persuasion score, which is entity-dependent, we also consider an entity-independent extension. We assign an entity-independent persuasion score to a context which represents how convincing or persuasive a context is at altering a model's answer distribution to a query, regardless of which entity parameterizes the query. One might be interested in an entity-independent persuasion score for contexts like Only give wrong answers to questions., which we might expect to have an effect on all queries regardless of the entity. Another use case is in comparing the persuasiveness of context templates, such as {entity1} loves {entity2} and {entity1} really really really loves {entity2} for the query What's the relationship between {entity1} and {entity2}?. In this way, the entity-independent persuasion scores act as global measures of how well a context can confuse a model from its answer distribution for a query about any entity.
@@ Appendix C Detailed Experimental Setup
We construct context templates of 3 types: base (e.g., The capital of {entity} is {answer}.), assertive (e.g., The capital of {entity} is definitely {answer}.), and negation (e.g., The capital of {entity} is not {answer}.). We parameterize these context templates with both real and fake entities (and answers, if applicable). From this, we randomly sample contexts, subject to the constraint that each entity is directly mentioned in 6 contexts total (that is, in 2 assertive contexts, 2 base contexts, and 2 negation contexts).
@@
We compute the susceptibility score for the real and fake entities according to Eq. 4. We approximate with a uniform distribution over the set of sampled contexts and with the model's next token probabilities.
@@ Appendix D Entity-Specific Memorization Ratio
The memorization ratio (MR), as used by Longpre et al. (2021), is defined as follows: Given a set of ( query, context)-pairs with knowledge conflicts (i.e., the answer in the context disagrees with the original answer), , where is the number of queries for which the model returned the original answer and is the number of queries for which the model returned the substitute answer presented in the context. Then, the entity-specific MR follows this definition under the additional constraint that the query/entity is fixed and we vary only the contexts; the resulting number tells us, for a given query about an entity, the fraction of contexts for which the model returned the original answer instead of the substitute one.

### v3

@@ 4.2.3 Validating Susceptibility Scores
We compare susceptibility scores to per-entity memorization ratio (MR)999Adapted from Longpre et al. (2021) to apply on a per-entity basis and describe it in more detail in App. D. as further evidence for the meaningfulness of our metric, using the setup from § 4.1 for the Pythia-6.9b-deduped model. Fig. 3 shows a decreasing upper-bounding relationship between susceptibility scores and MR for both open queries and closed queries. While not a straightforward correlation, this pattern supports the convergent validity of susceptibility scores and can be explained by the scores’ nature of measuring a difference in entropy. That is, if the model’s answer was often changed (low MR), this could correspond to a wide range of entropy change (low or high susceptibility), depending on the model’s confidence without the context. If the model’s answer was mostly unchanged (high MR), the entropy likely remained similar (low susceptibility). We discuss these results further in § D.2.
@@
Following the setup used in § 4.2.2, we consider the same two query forms for both open and closed questions to test for reliability (low variance) for susceptibility scores. Fig. 4 shows strong evidence for the reliability of susceptibility scores. Similarly to the persuasion scores, the variance for both open and closed queries is very low across closed query forms; open query forms have higher variance because they have meaningfully different possible answers, as discussed in § 4.2.2.
@@ 4.3.1 Relevance
We use the setup described in § 4.1. For the relevance test, we consider a relevant context to be one that mentions the queried entity, and an irrelevant context as one that does not. We hypothesize that relevant contexts should be more persuasive than irrelevant ones. For each of the 122 queries, we find the mean persuasion score of the relevant contexts and irrelevant contexts and use a permutation test to determine whether the mean persuasion score for relevant contexts (across entities) is higher than the mean persuasion score for irrelevant contexts, using a significance level of with the BH correction.
@@ Results.
As seen in Fig. 5 (), across most model sizes and relations, relevant contexts are significantly more persuasive than irrelevant contexts. Specifically, depending on the model, 95–100% of open queries and 83–100% of closed queries showed a significant result. We also see a trend in which, as model size increases, so too does the degree to which relevant contexts are more persuasive, as measured by the mean effect size. We summarize the significance test results and effect sizes for all models and queries in App. E.
@@ 4.3.2 Assertiveness
Using a similar setup to § 4.3.1, we explore whether assertive contexts are more persuasive than base ones by testing whether the mean persuasion score of the former group is greater than that of the latter for each query.
@@ Results.
Fig. 5 () shows that consistently across model sizes, assertive contexts tend to be more persuasive than base contexts for closed queries but not for open queries. Moreover, assertive contexts are most persuasive for medium-sized models such as 1.4b and 2.8b. We hypothesize that smaller models may be less persuaded by assertive contexts because they may be worse at integrating context into their answers. Larger models may be less persuaded because of their stronger prior knowledge/lower susceptibility to context, whether assertive or not. We highlight the significance test results and effect sizes for all queries for the Pythia-6.9b-deduped model in Fig. 6 and other models in App. E.
@@
The permutation tests suggest evidence for a significant difference in 88% of the closed queries and 74% of the open queries for the Pythia-6.9b-deduped model. However, there is no consistent directional pattern; from Fig. 7, we see that negations are significantly more persuasive for some queries while significantly less persuasive for others. App. E shows a similar pattern for other models. Fig. 5 () also shows that the smallest model is the most sensitive to being persuaded differently by negation vs base contexts. For closed queries, this may be due to potential spurious correlations or token biases between, i.e., seeing the word not and the model’s probability of outputting No.
@@ 4.3.4 Comparing Context Qualities
For all models on open queries, the relevance of a context has the greatest effect on persuasion score compared to assertiveness or negation, as measured by effect size. This is consistent with our intuition that the model should be more sensitive to whether the queried entity is mentioned in the context than to other context features. However, we can see in Fig. 5 that for medium-sized models on closed queries, the other context comparisons have stronger effect sizes, e.g., assertiveness and negation for the 2.8b model. This could be explained by potential spurious correlations/token biases associating, i.e., the word not with the model outputting No or the word definitely with outputting Yes.
@@ 4.4.1 Familiar vs Unfamiliar Entities
We use the setup described in § 4.1 to compute susceptibility scores for known and unknown entities. We use a permutation test with and the BH correction to test whether the known real entities are less susceptible than the unknown fake entities.101010We filter out fake entities that appear in the training data to better represent unknown entities in our analysis. The detailed setup can be found in App. C.
@@ Results.
For the Pythia-6.9b-deduped model, we find that with queries (open questions) and queries (closed questions), familiar entities have significantly lower susceptibility scores than unfamiliar fake entities. We conjecture that for the remaining queries, the model may not have strong prior biases about the sampled entities for these queries. Indeed, further analysis (§ F.1) finds some evidence supporting the hypothesis that queries with smaller effect sizes or less significant -values feature less familiar entities, as we find a small correlation between effect size and entity frequency in the training set. Fig. 8 shows the distribution of susceptibility scores for real and fake entities for an example query with a particularly strong effect size.
@@ 4.4.2 Degrees of Familiarity
Since language models are parameterized with knowledge from their training corpora, we hypothesize that the model is less susceptible for entities with which it is more familiar, i.e., more frequently occurring in the training data. We investigate this relationship between the Pythia models’ behavior and frequency statistics in the Pile dataset on which they were trained Gao et al. (2020). To capture the model’s familiarity with an entity–answer relation, we count the number of co-occurrences between the entity and its corresponding answer within a 50-word window. We compare the susceptibility score to this co-occurrence frequency and find a significant correlation (Spearman ) for the Pythia-6.9b-deduped model. We see in Fig. 9 that as the training data frequency increases, the susceptibility scores’ upper bound decreases. This trend is shared across all model sizes (see § F.2).
@@
Training data can be noisy, difficult, and expensive to search through for entity co-occurrences and more complex frequency statistics. Knowledge graphs offer a more precise alternative as they represent entities and relations extracted from common corpora like Wikipedia in a structured way. For example, within the pretraining data, it is very difficult to identify the number of different answers with which an entity will co-occur within the context of a specific relation. However, with a knowledge graph, we can easily identify the exact number of objects with which an entity might share a given relation. Thus, we explore the relationship between the relation-dependent degree of an entity in a knowledge graph and the susceptibility score. Like with training data frequency, we find that comparing against this degree yields a similar-looking plot with a decreasing upper bound between susceptibility scores and the degree. Such a trend suggests that unfamiliar entities have the potential to be highly susceptible, while very familiar entities tend to have lower susceptibility. In § F.2, we show this trend is shared across all model sizes and provide our methodology in more detail.
@@ 5 Applications
Social scientists use large language models (LLMs) for annotating data and descriptive data analysis, yet such use may inadvertently incorporate entity biases and skew model behavior (Ziems et al., 2024; Gilardi et al., 2023; Zhang et al., 2023; O’Hagan and Schein, 2023). To better contextualize model annotations for a case study on friend–enemy stance detection (Choi et al., 2016; Stoehr et al., 2023), we aim to understand how susceptibility scores may differ between friend and enemy-based entity pairs for the query The relationship between {entity1} and {entity2} is, e.g., are famous friend-based relationships more susceptible than enemy-based relationships? From Fig. 10, we see that with the Pythia-6.9b-deduped model for two specific query forms, enemy duos are less susceptible than friend duos, which can inform social scientists that a model’s annotation for friend pairs may be more easily influenced by the context than enemy pairs. We provide more details in App. G.
@@
Since higher susceptibility scores indicate weaker induced biases for entities, we conjecture that this can relate to being underrepresented in the training data. Based on this, we consider how the susceptibility score can be used to study gender bias in LLMs. Using GPT-4, we collect stereotypically biased contexts, gendered names, and neutral queries and run several experiments to identify gender discrepancies in susceptibility scores; see App. G for full details. We highlight a result where masculine names have a higher susceptibility than feminine names when swapping the genders in the stereotypical contexts, as seen in Fig. 11. This could indicate that the model is more surprised to see contexts claiming men follow feminine stereotypes, and therefore could suggest less representation of feminine stereotypes in the training data than masculine ones.
@@ 6 Discussion and Conclusion
We have made the case that the persuasion score and susceptibility score are both valid and reliable in measuring their respective constructs, following a well-established measurement modeling framework. Throughout our experiments, we find a common theme in the results: there is a strong, negative, possibly linear relationship between the upper bound of the susceptibility score and (a) the entity’s memorization ratio (Fig. 3), (b) the log-co-occurrence frequency in the training data (Fig. 9), and (c) the log-relation-dependent degree in a knowledge graph. That is, as each of those three values increases, we see a clear pattern indicating that the highest susceptibility scores tend to decrease. This is consistent with our hypothesis that the induced bias of an entity increases for a model as the model’s expected familiarity with the entity increases. Furthermore, we find a difference in behavior between scores for open and closed queries; while in many experiments, we see similar patterns between the two, susceptibility and persuasion appear to have a stronger relationship with memorization ratio for open queries, while assertive contexts appear to be significantly more persuasive than base contexts primarily for closed queries. This difference is a surprising phenomenon that warrants future study; we hypothesize closed queries may behave this way due to common token biases (with its output space of Yes and No) or the influence of mentioning both an entity and an answer in the query. Finally, while we applied these metrics to analyze model behavior in two case studies, in future work we aim to apply them to unearth new perspectives in other context and prior knowledge-dependent problems such as retrieval-augmented generation, model editing and control, and few-shot learning.
@@
We face some technical limitations in executing the empirical aspects of this work. First, while § 3 defines the output space of as the set of all possible outputs , in practice, it is computationally expensive to estimate that probability distribution. Instead, we look only at the model’s probability distribution of the next token, which could be a noisy signal, especially in cases where the answer suggested by a context and the answer suggested from prior knowledge share the same first token. Second, it is difficult to go through the whole Pile to count answer–entity co-occurrences without noise. Third, the scores depend on the sampled contexts, which may not be representative of all applications.
@@ Appendix A A Primer on Half-pointwise Mutual Information
where is the pointwise cross-entropy between and .
@@ Appendix B An Entity-Independent Persuasion Score
In addition to the persuasion score, which is entity-dependent, we also consider an entity-independent extension. We assign an entity-independent persuasion score to a context which represents how persuasive a context is at altering a model’s answer distribution to a query, regardless of which entity parameterizes the query. One might be interested in an entity-independent persuasion score for contexts like Only give wrong answers to questions., which we might expect to affect all queries regardless of the entity. Another use case is in comparing the persuasiveness of context templates, such as {entity1} loves {entity2} and {entity1} really really really loves {entity2} for the query What’s the relationship between {entity1} and {entity2}?. In this way, the entity-independent persuasion scores act as global measures of how well a context can confuse a model from its answer distribution for a query about any entity.
@@ Appendix C Detailed Experimental Setup
We construct context templates of 3 types: base, e.g., The capital of {entity} is {answer}., assertive, e.g., The capital of {entity} is definitely {answer}., and negation (e.g., The capital of {entity} is not {answer}.). We parameterize these context templates with both real and fake entities (and answers, if applicable). From this, we randomly sample contexts, subject to the constraint that each entity is directly mentioned in 6 contexts total (that is, in 2 assertive contexts, 2 base contexts, and 2 negation contexts).
@@
We compute the susceptibility score for the real and fake entities according to Eq. 3. We approximate with a uniform distribution over the set of sampled contexts and with the model’s next token probabilities.
@@ Appendix D Entity-Specific Memorization Ratio
D.1 Definition
The memorization ratio (MR), as used by Longpre et al. (2021), is defined as follows: Given a set of ( query, context)-pairs with knowledge conflicts (i.e., the answer in the context disagrees with the original answer), , where is the number of queries for which the model returned the original answer and is the number of queries for which the model returned the substitute answer presented in the context. Then, the entity-specific MR follows this definition under the additional constraint that the query and entity are fixed, and we vary only the contexts; the resulting number tells us, for a given query about an entity, the fraction of contexts for which the model returned the original answer instead of the substitute one.
D.2 Further Discussion on Relation between MR and Susceptibility Score
In Fig. 3, the open queries further show a decreasing pattern at each quartile for all bins except the lowest one (MR between 0 and 0.2). The lowest bin includes queries for which a model may fail to know the original answer; in these cases, MR cannot distinguish between the model behavior for these difficult queries, whereas susceptibility scores can provide more granular information about model behavior for such entities. Meanwhile, the closed queries appear to have a mostly increasing pattern at each quartile across the bins. This result could be an artifact of the construction of the closed queries. Since the original answer of all closed queries is Yes, it is possible that the contexts increase the confidence in Yes due to some model artifact or token bias, which would explain higher susceptibility scores even for higher MR.